# Centrally Coordinated Multi-Agent Reinforcement Learning for Power Grid Topology Control

### v1

@@ 2.2 Deep RL for Power Network Control
In L2RPN competitions, successful solutions typically combine expert rules, RL agents, and brute force simulations for action validation. Notably, RL agents based on Proximal Policy Optimization (PPO) [20], with reduced action spaces, have performed well. This approach was used in top-ranking competition entries (e.g. [21], further explored in studies [22, 23, 24]).
Several commonalities emerge among top-ranking approaches. Most successful methods implement intervention only during hazardous situations [25, 26, 23, 24, 27, 22]. In addition, many such approaches integrate learned modules with simulation-based action evaluation. For instance, this strategy was key to the success of the winning teams in both 2019 [26] and 2021 [28]. Notably, the second-place competitor in 2021 employed an advanced expert system, underscoring the value of incorporating domain knowledge from power systems [29].
@@ 4.1 Experimental Setup
Moreover, unplanned outages can be included. These outages involve temporarily disconnecting a powerline for a specified duration. These unplanned outages are initiated by an opponent, adding stochasticity and further increasing task complexity. This opponent introduces realistic challenges faced by power grid operators, ensuring the system’s stability is preserved even during unforeseen contingencies. It is designed to randomly disconnect a line from a predefined set of targets, with a cooldown period implemented to delay consecutive attacks. After each attack, the disconnected line is automatically restored. Table 1 outlines the parameters of these attacks.
@@ 4.2 Power System Feedback Control Elements
In this study, we introduce a new action space designed to ensure structural N-1 security. Structural N-1 security means that the grid remains operational and avoids an automatic blackout in the event of a single contingency, such as an unplanned powerline outage. We compare the new action space with the default action space which only excludes symmetrical counterparts from all possible substation configurations. Table 2 shows the size of each action space.
@@ 5.2 Multi-Agent Architecture
We divide the overall agent task into two levels of abstraction, following the approaches outlined in [23, 31]. One level consists of deciding the topological reconfiguration of individual regions (e.g. substations). The other level consists of selecting which region should be reconfigured. The agent selecting the region is called the coordinator. The reconfiguration of regions may be decided by multiple agents, called regional agents, each proposing a topological reconfiguration of its respective region. In our architecture the topological reconfiguration of each individual region is decided first, and the region to be reconfigured is selected second. Figure 2 illustrates on the left the high-level hierarchy used in Manczak et al. [23], van der Sar et al. [31] where the regional agents are called after the coordinator, and on the right the decision making order proposed in this study.
@@ 5.3 Overview of power system agents
Algorithm 1 Ordered CAPA Algorithm
2:  $current\_rho\leftarrow$ $max(rho)$ in $obs$
3:  if $P=None$ or $P$ is empty then
4:     Generate priority list $P$ using CAPA policy
5:  end if
6:  for each action $A$ do
7:     if $A\neq do\_nothing$ then
8:        Remove $A$ from $P$
9:        return $A$ and $P$
10:     end if
11:  end for
12:  $P\leftarrow None$
13:  return $do\_nothing$ and $P$
Table 3: The input for each module in the agent architecture.
Agent Name Input
Regional Agents
Greedy Observation
RL Observation
Coordinating Agents
CAPA Regional Actions & Observation
RL Regional Actions & Observation
Action-Value RL Regional Action-Value Pairs
& Observation
Value Softmax Regional Action-Value Pairs
@@ 5.4.2 Machine Learning Practices
Hyperparameter tuning is critical in RL research and applications because hyperparameters greatly affect the performance, stability, and efficiency of RL algorithms. For each architecture and environment, two grid searches are conducted to tune the parameters using Ray Tune [37]. The first is a coarse grid search to identify suitable learning rates, batch sizes, minibatch sizes, and the number of iterations. The second grid search is narrower, focusing on smaller variations in these parameters, as well as additional parameters such as the clipping range, value function coefficient range, and GAE lambda range. Appendix A details the full grid search settings for all experiments.
@@ 6 Results
Also in the most challenging 14-bus environment with opponent (lower-right panel in Fig. 4), Greedy-RL exhibits extraordinary sample efficiency and performance. The learning curves of Greedy-RL start at the performance level of the converged single RL baseline models. We note that for Greedy-RL the learning curves related to different seeds, shown in Figure 5 in Appendix B, exhibit bimodal behavior (in contrast to all other architectures). That is, one group of seeds remains at a lower performance level (but still higher than the level of the converged single RL baseline) whereas a second group of seeds exhibits steep learning curves outperforming both the single RL baseline and the rule-based benchmark approaches (i.e., the greedy agent and Greedy-CAPA) by a large margin. Hence, a mean curve (averaged across all seeds) is not a good representation of the model behavior such that in the lower-right panel of Fig. 4 we only show an average across high-performing seeds. More details on the training curves related to individual seeds can be found in App. B.
For completeness, Table 8 in Appendix C shows model evaluation on the test scenarios, which exhibits results consistent with the evaluation on the validation scenarios during training. However, the focus of this study is on the sample efficiency of agent architectures, and, hence, in this study the model evaluation during training (Fig. 4) is more relevant than final test results (Tab. 8 in App. C).
7 Discussion and Future Work
7.1 Discussion
@@
On the other hand, the Greedy-RL agent exhibits bi-modal behavior in the stochastic environment (Fig. 5 in App. B). One possible explanation is that not all contingency cases have equal value. It could be that only a smaller subset of the contingency cases pushes the agent to pick safer actions. This may have happened only on certain seeds. One possible explanation as to why we observe this only with Greedy-RL is its higher sample-efficiency. This would make the agent faster at picking up such patterns when (and if) they appear.
@@ 7.2 Future Work
Future research could build on this study by exploring several avenues for extending and refining the proposed methods. Testing the scalability and robustness of the approach in larger networks, such as a 118-bus network, could provide insights into its effectiveness in more complex environments. Moreover, future studies could test the impact of different strategies for decomposing the actions space [38].
@@
The results also suggest more concrete next steps for making the CCMA more scalable. The Greedy-RL architecture shows the most promise in sample efficiency. However, it is also the most computationally expensive agent. To mediate this, one could use imitation learning to train regional agents. These agents would try to imitate the greedy agent directly just much faster.
@@ Acknowledgments
Davide Grossi acknowledges support by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research (NWO)
Appendix A Hyperparameters
@@
A.1 5-bus network without opponent
@@
A.2 5-bus network with opponent
Grid search was performed over the following parameters:
  • Learning rate (lr): 0.001, 0.0001, 0.00001, 0.000001
  • SGD minibatch size (mbs): 32, 64, 128, 256
  • Train batch size (bs): 128, 256, 512, 1024
  • Number of SGD iterations (it): 5, 10
@@
A.3 14-bus network without opponent
@@
A.4 14-bus network with opponent
@@
Appendix B Individual Seeds for 14-Bus Network With Opponent
@@
Appendix C Results of test scenarios

### v2

@@ 2.2 Deep RL for Power Network Control
In L2RPN competitions, successful solutions typically combine expert rules, RL agents, and brute force simulations for action validation. Notably, RL agents based on Proximal Policy Optimization (PPO) [35], with reduced action spaces, have performed well. This approach was used in top-ranking competition entries (e.g. [28], further explored in studies [17, 26, 3, 16]).
Several commonalities emerge among top-ranking approaches. Most successful methods implement intervention only during hazardous situations [42, 15, 26, 3, 9, 17, 23]. In addition, many such approaches integrate learned modules with simulation-based action evaluation. For instance, this strategy was key to the success of the winning teams in both 2019 [15] and 2021 [32]. Notably, the second-place competitor in 2021 employed an advanced expert system, underscoring the value of incorporating domain knowledge from power systems [27].
@@ 4.1 Experimental Setup
Moreover, unplanned outages can be included. These outages involve temporarily disconnecting a powerline for a specified duration. These unplanned outages are initiated by an opponent, adding stochasticity and further increasing task complexity. This opponent introduces realistic challenges faced by power grid operators, ensuring the system’s stability is preserved even during unforeseen contingencies. The opponent is implemented as Grid2Op’s RandomLineOpponent [6]. It is designed to randomly disconnect a line from a predefined set of targets, with a cooldown period implemented to delay consecutive attacks, identical to Manczak et al. [26]. After each attack, the disconnected line is automatically restored. Table 1 outlines the parameters of these attacks.
@@ 4.2 Power System Feedback Control Elements
In this study, we introduce a new action space designed to ensure structural N-1 security. Structural N-1 security means that the grid remains operational and avoids an automatic blackout in the event of a single contingency, such as an unplanned powerline outage. We compare the new action space with the default action space which only excludes symmetrical counterparts from all possible substation configurations.
Table 2 shows the size of each action space.
@@ 5.2 Multi-Agent Architecture
We divide the overall agent task into two levels of abstraction, following the approaches outlined in Manczak et al. [26], van der Sar et al. [39]. One level consists of deciding the topological reconfiguration of individual regions (e.g. substations). The other level consists of selecting which region should be reconfigured. The agent selecting the region is called the coordinator. The reconfiguration of regions may be decided by multiple agents, called regional agents, each proposing a topological reconfiguration of its respective region. In our architecture the topological reconfiguration of each individual region is decided first, and the region to be reconfigured is selected second. Figure 2 illustrates on the left the high-level hierarchy used in Manczak et al. [26], van der Sar et al. [39] where the regional agents are called after the coordinator, and on the right the decision making order proposed in this study.
@@
Algorithm 1 Ordered CAPA Algorithm
2:  $current\_rho\leftarrow$ $max(rho)$ in $obs$
3:  if $P=None$ or $P$ is empty then
4:     Generate priority list $P$ using CAPA policy
5:  end if
6:  for each action $A$ do
7:     if $A\neq do\_nothing$ then
8:        Remove $A$ from $P$
9:        return $A$ and $P$
10:     end if
11:  end for
12:  $P\leftarrow None$
13:  return $do\_nothing$ and $P$
@@
Table 4 in Appendix A highlights the key differences between the proposed CCMA architecture and other state-of-the-art architectures in power grid topology control.
@@ 5.3 Overview of power system agents
Table 3: The input for each module in the agent architecture.
Agent Name Input
Regional Agents
Greedy Observation
RL Observation
Coordinating Agents
CAPA Regional Actions & Observation
RL Regional Actions & Observation
Action-Value RL Regional Action-Value Pairs & Observation
Value Softmax Regional Action-Value Pairs
@@ 5.4.2 Machine Learning Practices
Hyperparameter tuning is critical in RL research and applications because hyperparameters greatly affect the performance, stability, and efficiency of RL algorithms. For each architecture and environment, two grid searches are conducted to tune the parameters using Ray Tune [20]. The first is a coarse grid search to identify suitable learning rates, batch sizes, minibatch sizes, and the number of iterations. The second grid search is narrower, focusing on smaller variations in these parameters, as well as additional parameters such as the clipping range, value function coefficient range, and GAE lambda range. Appendix B details the full grid search settings for all experiments.
It should be noted that to validate the stability and reliability of the proposed method, all architectures are trained independently per network using ten different random seeds. The stability of the approach is illustrated explicitly in Figure 4, where the variability across multiple seeds is represented by the standard deviation, indicating the robustness and convergence behavior under diverse initialization conditions.
@@ 6 Results
Also in the most challenging 14-bus environment with opponent (lower-right panel in Fig. 4), Greedy-RL exhibits extraordinary sample efficiency and performance. The learning curves of Greedy-RL start at the performance level of the converged single RL baseline models. We note that for Greedy-RL the learning curves related to different seeds, shown in Figure 5 in Appendix C, exhibit bimodal behavior (in contrast to all other architectures). That is, one group of seeds remains at a lower performance level (but still higher than the level of the converged single RL baseline) whereas a second group of seeds exhibits steep learning curves outperforming both the single RL baseline and the rule-based benchmark approaches (i.e., the greedy agent and Greedy-CAPA) by a large margin. Hence, a mean curve (averaged across all seeds) is not a good representation of the model behavior such that in the lower-right panel of Fig. 4 we only show an average across high-performing seeds. More details on the training curves related to individual seeds can be found in App. C.
For completeness, Table 9 in Appendix D shows model evaluation on the test scenarios, which exhibits results consistent with the evaluation on the validation scenarios during training. However, the focus of this study is on the sample efficiency of agent architectures, and, hence, in this study the model evaluation during training (Fig. 4) is more relevant than final test results (Tab. 9 in App. D).
7 Discussion
@@
On the other hand, the Greedy-RL agent exhibits bi-modal behavior in the stochastic environment (Fig. 5 in App. C). One possible explanation is that not all contingency cases have equal value. It could be that only a smaller subset of the contingency cases pushes the agent to pick safer actions. This may have happened only on certain seeds. One possible explanation as to why we observe this only with Greedy-RL is its higher sample-efficiency. This would make the agent faster at picking up such patterns when (and if) they appear.
@@
As a final note, Fig. 4 illustrates the stability of the proposed method by illustrating the standard deviation over ten different random seeds and multiple validation scenarios used during training. It stands out that some architectures exhibit considerably lower variability across validation scenarios and seeds, as reflected by smaller standard deviations, indicating more stable and robust convergence behavior. In contrast, other architectures show higher variability, suggesting that their performance is more sensitive to initialization.
8 Conclusion and Future Work
@@ 8.1 Future Work
Future research could build on this study by exploring several avenues for extending and refining the proposed methods. Testing the scalability and robustness of the approach in larger networks, such as the 36-bus [27] and 118-bus networks [36], could provide insights into its effectiveness in more complex environments. To assess scalability, we conducted a preliminary evaluation on a 36-bus network (66,810 possible topological actions) and find promising outcomes. The results are shown in Appendix E. Moreover, future studies could test the impact of different strategies for decomposing the actions space [24].
@@
The results also suggest more concrete next steps for making the CCMA more scalable. The Greedy-RL architecture shows the most promise in sample efficiency. However, it is also the most computationally expensive agent as detailed in Appendix F. To mediate this, one could use imitation learning to train regional agents. These agents would try to imitate the greedy agent directly just much faster.
@@ Acknowledgments
Davide Grossi acknowledges support by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research (NWO).
Appendix A Comparative Analysis of Power Grid Control Architectures
Table 4 compares the proposed CCMA architecture and other state-of-the-art architectures, focusing on agent architecture, action space representation and action space reduction.
Table 4: Comparison of RL-based Architectures for Power Grid Topology Control.
Paper Agent Architecture Type of Action Space for RL Agent Action Space Reduction
CCMA, proposed in this study Hierarchical multi-agent architecture that separates decision layers between, first, regional agents and, second, a coordinator. Distributed proposals over multiple regional agents yield a diverse range of candidate actions. Directly addresses the combinatorial challenge by decomposing the action space, so the problem scales with the sum of sub-problem sizes rather than their product. RL coordinator learns on higher-level action space consisting of the set of controllable substations. This action space is naturally much smaller than the primitive action space. No arbitrary action space reduction is necessary. Only power system principles are used to reduce the action space, without relying on heuristics or manual tuning. This also avoids introducing bias through greedy action selection.
Liu et al. [23] Hierarchical agent architecture similar to Manczak et al. [26], as discussed in Sec. 5.2 in this study. That is, first a coordinator specifies a substation, and subsequently a single regional agent specifies the substation configuration. As in CCMA, the coordinator learns on higher-level action space consisting of the set of controllable substations. However, the coordinator does not receive the output of the regional agent as input, as discussed in Sec. 5.2 in this study. Similar to CCMA, that is, no action space reduction but action space decomposition. However, no action space decomposition on the subsation level.
Yoon et al. [42] Hierarchical approach consisting of, first, a RL agent choosing a goal topology, and, second, a rule-based agent to specify the corresponding sequence of primitive substation changes. Moreover, an afterstate representation is used. RL agent learns on entire topology action space which, hence, needs to be drastically reduced. The action space consists of goal topologies, which suffers even more from combinatorial explosion than substation reconfiguration actions. How the action space is reduced is not discussed in the paper, but the code includes an arbitrary cut-off range of maximum possible actions per substation.
Lehna et al. [17] Employs a teacher-tutor-junior-senior framework. Essentially, this is a single RL approach with a warm-start based on greedy expert data. RL agent needs to learn on the primitive action space which, hence, needs to be drastically reduced. Uses brute-force greedy search to create a frequency distribution of the entire action space. Subsequently, the action space is then drastically pruned based on frequency distribution using an arbitrary cut-point.
Lehna et al. [16] Extending Lehna et al. [17] by adding a rule-based application of so-called target topologies in a low loading regime. Same as Lehna et al. [17], that is, RL agent needs to learn on the primitive action space which, hence, needs to be drastically reduced. Similar to Lehna et al. [17]. Reduced set of target topologies is based greedy agent statistics. The reduced set of primitive actions is simply taken from another agent by Dorfer et al. [9].
Chauhan et al. [3] Essentially a single RL approach supported by a number of domain specific heuristics similar to Lehna et al. [17], see also Sec. 4.1.1 in this study. Same as Lehna et al. [17], that is, RL agent needs to learn on the primitive action space which, hence, needs to be drastically reduced. Similar to Lehna et al. [17]. Through extensive brute-force simulation, the action space is drastically reduced; the actual action space reduction procedure is not explained.
Dorfer et al. [9] Essentially a single RL approach using Monte Carlo Tree Search. Same as Lehna et al. [17], that is, RL agent needs to learn on the primitive action space which, hence, needs to be drastically reduced. Similar to Lehna et al. [17]. Through extensive brute-force simulation, the action space is drastically reduced by using the top most frequent actions.
Appendix B Hyperparameters
@@
B.1 5-bus network without opponent
@@
B.2 5-bus network with opponent
Grid search was performed over the following parameters:
  • Learning rate (lr): 0.001, 0.0001, 0.00001, 0.000001
  • SGD minibatch size (mbs): 32, 64, 128, 256
  • Train batch size (bs): 128, 256, 512, 1024
  • Number of SGD iterations (it): 5, 10
@@
B.3 14-bus network without opponent
@@
B.4 14-bus network with opponent
@@
Appendix C Individual Seeds for 14-Bus Network With Opponent
@@
Appendix D Results of test scenarios
@@
Appendix E Preliminary 36-bus Experiment
To explore scalability on the 36-bus network (l2rpn_icaps_2021_large), we experimented with the Greedy, Single RL and Greedy-RL agents.
First, we deployed a more efficient greedy agent that evaluates all non-hub substations (fewer than 1,000 actions) and executes the first move whose simulated peak line load remains below the safety threshold or decreases by at least 5%, thus avoiding full combinatorial search over hub actions. Only if no non-hub action qualifies, the agent continuously samples from the 65,121 hub actions until a good action is found. If still none meet the criteria, it selects the action minimizing the maximum line loading.
Next, we built a reduced action space by collecting every hub action chosen more than once by this greedy agent (while retaining all non-hub actions). We used this action space to train four and two sets of hyperparameters of both the Greedy-RL and Single-RL agents, respectively. Based on App. B, we evaluated the parameter setting specified in Table 10.
Table 10: Hyperparameter configurations for Greedy-RL and Single-RL on the 36-bus network without opponent.
Agent lr mbs bs it cp vfc vfl $\lambda$
Greedy-RL 0.0001 512 3072 10 0.1 20 0.95 0.95
Greedy-RL 0.00001 512 3072 10 0.1 20 0.95 0.95
Greedy-RL 0.00005 1024 3072 5 0.2 20 0.95 0.95
Greedy-RL 0.00005 1024 3072 5 0.2 20 0.95 0.95
Single-RL 0.00005 512 3072 10 0.1 20 0.95 0.95
Single-RL 0.00005 512 3072 10 0.1 20 0.95 0.95
Figure 6 shows the resulting training (a) and validation (b) curves. The results show that both the Single RL and Greedy-RL suffer from overfitting, suggesting that further hyperparameter tuning is needed. While the training curves consistently increase (see Fig. 6(a)), the validation curves show large variations (see Fig. 6(b)). It should be noted that when an agent is updated, its buffer will contain experiences from only a fraction of the training scenarios. One likely explanation for this overfitting (and the subsequent recovery) is that out of distribution training scenarios are picked for training (at around 200k environmental interactions in Fig. 6(b)). This most likely results in the agent trying to perform well in these peculiar scenarios, effectively “forgetting" some of the behaviors useful in the validation set.
(a) The rolling mean (window = 50) number of timesteps survived on training scenarios for the 36-bus network without opponent. As described in Sec. 4.1.1, the agent is, whenever feasible, trained on two-day segments (576 timesteps) extracted from the original training scenarios. When nighttime instabilities prevent a full two-day run, these segments may occasionally be one day shorter or longer. Across all training scenarios, the average segment length is 588.1 timesteps.
(b) The mean number of timesteps survived on all 296 validation scenarios throughout training for the 36-bus network without opponent.
Figure 6: Performance of the Greedy, Greedy-RL and Single RL agents on the 36-bus network without opponent over time: (a) training, (b) validation.
Regardless, the plots still highlight the value of Greedy-RL in terms of performance. The best validation performance achieved by Greedy-RL far surpasses that of both the Single-RL and the rule-based Greedy agent. These findings confirm our smaller-grid conclusions while underscoring the need for a full hyperparameter grid search. Future work will also explore splitting the hub agent into multiple regional agents, allowing the full range of actions to be used again.
Appendix F Computational Time
Table 11 presents the average computational time per environment interaction during training for various agents on the 14-bus network with an opponent, using a single CPU per worker. The Single RL agent demonstrates the lowest computational time, followed closely by the RL-RL and RL-Action Value RL agents. In contrast, the Greedy-RL agent exhibits significantly higher computational time, approximately an order of magnitude greater than the others.
Table 11: Compute Time Per Timestep While Training 14-Bus Network With Opponent
Agent Time per Environment Interaction (s)
Single RL 0.307
RL-RL 0.492
RL-Action Value RL 0.468
Greedy-RL 5.089